# Light and heavy $\Lambda$ hyperclusters in nuclear matter with RMF models

### v1

@@ Abstract
In the framework of RMF models, we investigate the properties of light and heavy $\Lambda$ hyperclusters emersed in nuclear matter at various densities $n_{\mathrm{gas}}$ and proton fractions $Y_{p}$. In particular, the (hyper)clusters are fixed by solving the Dirac equations imposing the Dirichlet-Neumann boundary condition, while the nuclear matter take constant densities and is treated with Thomas-Fermi approximation. The binding energies of (hyper)clusters decrease with the density of nuclear matter $n_{\mathrm{gas}}$, which eventually become unbound and melt in the presence of nuclear medium, i.e., Mott transition. For light clusters with proton numbers $N_{p}<4$, with the addition of $\Lambda$ hyperons, the binding energies per baryon for $\Lambda$ hyperclusters become smaller and decrease faster with $n_{\mathrm{gas}}$ due to the weaker $N$-$\Lambda$ attraction. For heavy clusters with $N_{p}\geq 4$, on the contrary, the addition of $\Lambda$ hyperons increases the stability of (hyper)clusters so that the Mott transition density becomes larger as nucleons occupying higher energy states while $\Lambda$ hyperons remain in the $1s_{1/2}$ orbital. The isovector effects on (hyper)clusters in nuclear medium are also identified, where the binding energies for (hyper)clusters with $N_{p}>N_{n}$ ($N_{p}<N_{n}$) increase (decrease) with $Y_{p}$. For those predicted by nonlinear relativistic density functionals, light (hyper)clusters are destabilized drastically as $n_{\mathrm{gas}}$ increases, while the binding energies of heavier (hyper)clusters vary smoothly with $n_{\mathrm{gas}}$. By fitting the binding energy shifts to an analytical formula, the corresponding coefficients describing the in-medium properties of various (hyper)clusters are fixed, which should be useful to understand the evolutions of (hyper)clusters in both heavy-ion collisions and neutron stars.
@@ I Introduction
Here $M({}^{A-1}Z)$, $M_{\Lambda}$, and $M({}^{A}_{\Lambda}Z)$ are the masses of nucleus ${}^{A-1}Z$, $\Lambda$ hyperon, and $\Lambda$ hypernuclues ${}^{A}_{\Lambda}Z$. Meanwhile, double $\Lambda$ hypernuclei have also been observed, such as ${}^{6}_{\Lambda\Lambda}$e, ${}^{10,11,12}_{\Lambda\Lambda}$e, and ${}^{12,13}_{\Lambda\Lambda}$oki et al. (2009); Ahn et al. (2013). The $\Lambda\Lambda$ binding energies $B_{\Lambda\Lambda}$ can then be obtained, from which the $\Lambda$$\Lambda$ interaction energy $\Delta B_{\Lambda\Lambda}$ is fixed via
@@
In heavy-ion collisions (with beam energy approximately larger than 1.58 GeV per nucleon), hyperons and hypernuclei can also be produced under the extreme conditions of high energy densities Chen et al. (2025). The existence of $\rm{{}^{3}_{\Lambda}H}$ in heavy-ion reactions were observed The STAR Collaboration (2020); ALICE Collaboration (2023), while heavier (multi-strangeness) hypernulcei are expected to be produced as well Cheng and Feng (2022). Nevertheless, their detailed production mechanism is still unclear, where two different formalisms were adopted to describe the production of these states, i.e., the statistical thermal method Andronic et al. (2018) and the coalescence model She et al. (2021). Due to the differences in $N$-$\Lambda$ and $N$-$N$ interactions, the production of hypernuclei has distinctive features in heavy-ion collisions compared with the corresponding normal nuclei of the same mass number Botta et al. (2012). Meanwhile, for heavy-ion collisions with beam energy close to hyperon threshold, higher-density nuclear matter is produced and the $N$-$N$ and $N$-$\Lambda$ interactions are expected to impact the hyperon dynamics, e.g., collective flows, transverse momentum spectra, and rapidity distribution Yong (2025, 2024); Yong et al. (2025); Feng (2024); Wei et al. (2024). As hypernuclei are produced in a hot and/or dense environment in heavy-ion collisions, the in-medium binding energy shift of hyperclusters maybe important for understanding hypernuclei production mechanism, particularly at beam energies around the hyperon threshold.
@@
In such cases, in this work we investigate the clustering phenomenon of $\Lambda$ hypernuclei in nuclear medium. In our previous work we have proposed a unified treatment for investigating the clustering phenomenon of normal nuclei inside a spherical Wigner-Seitz cell employing RMF models Xia (2025), which will be extended to the hyperonic sector here. The clusters are fixed by solving the Dirac equations imposing the Dirichlet-Neumann boundary condition Negele and Vautherin (1973), while the nuclear medium is treated with Thomas-Fermi approximation and assume constant densities. By calibrating the coupling constants $g_{\sigma\Lambda}$ and $g_{\omega\Lambda}$ according to the experimental separation energies of $\Lambda$ hyperons $B_{\Lambda}$ for single $\Lambda$ hypernuclei Hashimoto and Tamura (2006); Gal et al. (2016), the properties of $\Lambda$ hypernuclei in vacuum and in nuclear matter can be fixed. With the inclusion of $\Lambda$ hyperons, the binding energy of a $\Lambda$ hypernucleus ${}^{A}_{\Lambda}Z$ is larger than that of the core nucleus ${}^{A-1}Z$ with the additional $\Lambda$ binding energy $B_{\Lambda}$. In such cases, $\Lambda$ hyperclusters become more stable than normal clusters in nuclear medium, where the binding energies generally decline slower for $\Lambda$ hyperclusters with $A\gtrsim 8$ as the density of nuclear medium increases.
@@ II Theoretical framework
The Lagrangian density for RMF models can be obtained with
@@
where $G_{n\kappa}(r)/r$ and $F_{n\kappa}(r)/r$ are the upper and lower components for the radial wave functions, $Y_{jm}^{l}(\theta,\phi)$ is the spinor spherical harmonics, $\kappa$ is connected to the angular momenta $(l,j)$ via $\kappa=(-1)^{j+l+1/2}(j+1/2)$. The Dirac equation is then reduce into
@@
where reflective boundary conditions are imposed at $r=0$ and the edge of Wigner-Seitz cell with $r=R_{\mathrm{W}}$ Xia et al. (2021). Note that the global charge neutrality is imposed by adding negatively charged electron background $n_{e}$ so that $A(R_{\mathrm{W}})=\nabla A(R_{\mathrm{W}})=0$. The source currents are comprised of contributions from both the nuclear medium (gas) and cluster, i.e.,
@@
The scalar and vector densities of clusters (with given particle numbers $N_{i}$) are fixed according to the radial wave functions, i.e.,
@@
Note that the mass difference between nucleons and hyperons are neglected for estimating the center-of-mass correction to the proton radius.
@@
with $g(x)=x(2x^{2}+1)\sqrt{x^{2}+1}-\mathrm{arcsh}(x)$. The center-of-mass energy needs to be subtracted from Eq. (29) and can be fixed by Bender et al. (2000)
@@
where $S_{i,\mathrm{gas}}$ and $V_{i,\mathrm{gas}}$ are the scalar and vector mean-field potentials of baryon $i$ in the uniform nuclear medium.
@@ III Results and discussions
In this work we adopt the relativistic density functionals DD-LZ1 Wei et al. (2020), TM1 Sugahara and Toki (1994b), and PK1 Long et al. (2004), where the corresponding saturation properties of nuclear matter are listed in Table. 1. Note that the functionals PK1 and DD-LZ1 were widely adopted in our previous investigations on the properties of (hyper)nuclei Liu et al. (2018); Ding et al. (2023), while that of TM1 was extensively used for supernova simulations, i.e., the Shen EoSs Shen et al. (2011).
@@ III.1 $\Lambda$ hypernuclei
We then investigate the properties of $\Lambda$ hypernuclei based on these effective $N$-$N$ interactions in the absence of pairing interactions. For the $N$-$\Lambda$ interactions, we take the depth of $\Lambda$ mean field potential in symmetric nuclear matter (SYM) at saturation density to be
@@
which is expected to reproduce the single-$\Lambda$ binding energies irrespective of the particular values of $g_{\sigma\Lambda}$ and $g_{\omega\Lambda}$ Wang et al. (2013); Sun et al. (2018); Rong et al. (2021). Consider the two-solar-mass constraint on neutron stars’ masses Sun et al. (2018, 2019), we take $g_{\sigma\Lambda}=g_{\sigma N}$ and then fix $g_{\omega\Lambda}$ according to Eq. (35), which gives $g_{\omega\Lambda}\approx 1.14g_{\omega N}$.
@@
Figure 1: Binding energies of $\Lambda$ hyperon $B_{\Lambda}$ in the $1s_{1/2}$ state of hypernuclei predicted by the relativistic density functionals DD-LZ1 Wei et al. (2020), TM1 Sugahara and Toki (1994b), and PK1 Long et al. (2004). The corresponding experimental values are indicated by the black squares for comparison Hashimoto and Tamura (2006); Gal et al. (2016). The binding energy per nucleon $B_{\mathrm{core}}/(A-1)$ for the nuclear cores are presented as well.
@@
The binding energies of $\Lambda$ hyperon in the $1s_{1/2}$ states of $\Lambda$ hypernuclei (${}^{4}_{\Lambda}$, ${}^{4}_{\Lambda}$e, ${}^{7}_{\Lambda}$e, ${}^{7}_{\Lambda}$e, ${}^{8}_{\Lambda}$i, ${}^{10}_{\Lambda}$e, ${}^{10}_{\Lambda}$, ${}^{12}_{\Lambda}$, ${}^{12}_{\Lambda}$, ${}^{13}_{\Lambda}$, ${}^{13}_{\Lambda}$, ${}^{16}_{\Lambda}$, ${}^{16}_{\Lambda}$, ${}^{28}_{\Lambda}$i, ${}^{32}_{\Lambda}$, ${}^{40}_{\Lambda}$a, ${}^{51}_{\Lambda}$, ${}^{52}_{\Lambda}$, ${}^{89}_{\Lambda}$, ${}^{139}_{\Lambda}$a, and ${}^{208}_{\Lambda}$b) are then estimated with Eq. (1). The obtained results are presented in Fig. 1, while the corresponding experimental values are indicated as well for comparison Hashimoto and Tamura (2006); Gal et al. (2016). Evidently, the relativistic density functionals adopted in this work generally well reproduce the experimental binding energies of $\Lambda$ hyperon in $\Lambda$ hypernuclei in a large mass range of $A=4\sim 208$, where the corresponding root-mean-square deviations of $B_{\Lambda}$ from the experimental values are $\Delta=1.30$, 0.74, and 0.86 MeV when the functionals DD-LZ1 Wei et al. (2020), TM1 Sugahara and Toki (1994b), and PK1 Long et al. (2004) are adopted for the effective $NN$ interactions. The binding energy per nucleon $B_{\mathrm{core}}/(A-1)$ for the nuclear cores are presented as well. For light nuclei such as 3H and 3He, their binding energies per nucleon are larger than the binding energies of $\Lambda$ hyperon for ${}^{4}_{\Lambda}$ and ${}^{4}_{\Lambda}$e, which is mainly due to the weaker $N$-$\Lambda$ interaction than that of $N$-$N$ interaction. As mass number ($A\gtrsim 8$) increases, nucleons eventually occupy single-particle states of larger energies, while $\Lambda$ hyperons still occupy the $1s_{1/2}$ state, so that the binding energies of $\Lambda$ hyperon eventually surpass that of nucleons.
@@
Table 2: Binding energies $B$ and charge radii $R_{\mathrm{ch}}$ for finite nuclei predicted by the relativistic density functional DD-LZ1 Wei et al. (2020). By adding one (two) $\Lambda$ hyperon(s) to these nuclei, the hyperon separation energies $B_{\Lambda}$ ($B_{\Lambda\Lambda}$) of $\Lambda$ ($\Lambda$$\Lambda$) hypernuclei as well as their charge radii $R_{\mathrm{ch}}^{\Lambda}$ ($R_{\mathrm{ch}}^{\Lambda\Lambda}$) are obtained and listed here.
Core $B/A$ $R_{\mathrm{ch}}$ $B_{\Lambda}$ $R_{\mathrm{ch}}^{\Lambda}$ $B_{\Lambda\Lambda}/2$ $R_{\mathrm{ch}}^{\Lambda\Lambda}$
MeV fm MeV fm MeV fm
3H 3.747 2.090 2.00 2.164 1.923 2.187
3He 3.419 2.337 2.06 2.407 1.972 2.414
4He 6.832 1.988 2.71 2.030 2.644 2.020
8Be 5.220 2.464 7.50 2.357 6.969 2.292
12C 7.663 2.354 11.48 2.313 11.05 2.292
16O 8.028 2.671 10.98 2.637 10.68 2.616
40Ca 8.610 3.440 16.64 3.415 16.38 3.397
48Ca 8.659 3.440 18.93 3.416 18.67 3.399
208Pb 7.871 5.485 25.49 5.468 25.35 5.452
We then investigate the properties of various $\Lambda$ hypernuclei and $\Lambda$$\Lambda$ hypernuclei that may be produced in heavy-ion collisions as well as in violent astrophysical processes such as supernova explosions and binary neutron star mergers. As an example, we employ the relativistic density functional DD-LZ1 Wei et al. (2020) for the effective $N$-$N$ interactions and $g_{\omega\Lambda}=1.143g_{\omega N}$ ($g_{\sigma\Lambda}=g_{\sigma N}$) for the $\Lambda$-$N$ interactions. The obtained binding energy $B$ and charge radii $R_{\mathrm{ch}}$ of finite nuclei are then summarized in Table 2, which are generally consistent with the experimental data Angeli and Marinova (2013); Huang et al. (2021); Wang et al. (2021) except for the lightest nuclei at $A=3$ with $\Delta B/A\approx 1$ MeV and $\Delta R_{\mathrm{ch}}\approx 0.3$ fm Xia (2025). This should be improved in our future works by properly incorporating various corrections to the mean field approximation employed here, e.g., the inclusion of proton-neutron pairing Bally et al. (2025). By adding one or two $\Lambda$s to these nuclei, we can further obtain the binding energies $B_{\Lambda}$ and $B_{\Lambda\Lambda}$ for single $\Lambda$ and double $\Lambda$s in $\Lambda$ hypernuclei and $\Lambda$$\Lambda$ hypernuclei, while the corresponding charge radii $R_{\mathrm{ch}}^{\Lambda}$ and $R_{\mathrm{ch}}^{\Lambda\Lambda}$ are fixed as well. As indicated in Table 2, due to the attractive $\Lambda$-$N$ interaction, the binding energy of $\Lambda$ hyperon is always positive and increases with the mass number $A$, which are expected to reach $B_{\Lambda}=30$ MeV as $A\rightarrow\infty$ according to Eq. (35). Except for the light nuclei with $A=3$ or 4, the charge radii decrease with the inclusion of $\Lambda$ hyperons, i.e., the core becomes more compact due to the impurity effects of $\Lambda$ hyperons. Based on the binding energies $B_{\Lambda}$ and $B_{\Lambda\Lambda}$, the $\Lambda$-$\Lambda$ interaction energy $\Delta B_{\Lambda\Lambda}$ can be obtained, which are negative and correspond to a repulsive $\Lambda$-$\Lambda$ interaction. This contradicts the experimental constraints as $\Lambda$-$\Lambda$ interaction was found to be attractive in ${}^{6}_{\Lambda\Lambda}$e Takahashi et al. (2001); Ahn et al. (2013).
@@
Figure 2: Density profiles of 4He, ${}^{5}_{\Lambda}$e, and ${}^{6}_{\Lambda\Lambda}$e in vacuum, where the relativistic density functional DD-LZ1 Wei et al. (2020) is employed for the effective $N$-$N$ interactions and $g_{\omega\Lambda}=1.143g_{\omega N}$ ($g_{\sigma\Lambda}=g_{\sigma N}$) for the $\Lambda$-$N$ interactions.

### v2

# Light and heavy $\Lambda$ hyperclusters in nuclear matter with relativistic-mean-field models

@@ Abstract
In the framework of relativistic-mean-field (RMF) models, we investigate the properties of light and heavy $\Lambda$ hyperclusters emersed in nuclear matter at various densities $n_{\mathrm{gas}}$ and proton fractions $Y_{p}$. In particular, the (hyper)clusters are fixed by solving the Dirac equations imposing the Dirichlet-Neumann boundary condition, while the nuclear matter take constant densities and is treated with Thomas-Fermi approximation. The binding energies of (hyper)clusters decrease with the density of nuclear matter $n_{\mathrm{gas}}$, which eventually become unbound and melt in the presence of nuclear medium, i.e., Mott transition. For light clusters with proton numbers $N_{p}<4$, with the addition of $\Lambda$ hyperons, the binding energies per baryon for $\Lambda$ hyperclusters become smaller and decrease faster with $n_{\mathrm{gas}}$ due to the weaker $N$-$\Lambda$ attraction. For heavy clusters with $N_{p}\geq 4$, on the contrary, the addition of $\Lambda$ hyperons increases the stability of (hyper)clusters so that the Mott transition density becomes larger as nucleons occupying higher energy states while $\Lambda$ hyperons remain in the $1s_{1/2}$ orbital. The isovector effects on (hyper)clusters in nuclear medium are also identified, where the binding energies for (hyper)clusters with $N_{p}>N_{n}$ ($N_{p}<N_{n}$) increase (decrease) with $Y_{p}$. For those predicted by nonlinear relativistic density functionals, light (hyper)clusters are destabilized drastically as $n_{\mathrm{gas}}$ increases, while the binding energies of heavier (hyper)clusters vary smoothly with $n_{\mathrm{gas}}$. The binding energy shifts of various (hyper)clusters due to the impact of nuclear medium are fitted to an analytical formula, which could be employed to examine the evolutions of (hyper)clusters in both heavy-ion collisions and neutron stars.
@@ I Introduction
Here $M({}^{A-1}Z)$, $M_{\Lambda}$, and $M({}^{A}_{\Lambda}Z)$ are the masses of nucleus $^{A-1}Z$, $\Lambda$ hyperon, and $\Lambda$ hypernucleus $^{A}_{\Lambda}Z$. Meanwhile, double $\Lambda$ hypernuclei have also been observed, such as $^{6}_{\Lambda\Lambda}$He, $^{10,11,12}_{\Lambda\Lambda}$Be, and $^{12,13}_{\Lambda\Lambda}$Aoki et al. (2009); Ahn et al. (2013). The $\Lambda\Lambda$ binding energies $B_{\Lambda\Lambda}$ can then be obtained, from which the $\Lambda$$\Lambda$ interaction energy $\Delta B_{\Lambda\Lambda}$ is fixed via
@@
In heavy-ion collisions (with beam energy approximately larger than 1.58 GeV per nucleon), hyperons and hypernuclei can also be produced under the extreme conditions of high energy densities Chen et al. (2025). The existence of $\rm{{}^{3}_{\Lambda}H}$ in heavy-ion reactions were observed The STAR Collaboration (2020); ALICE Collaboration (2023), while heavier (multi-strangeness) hypernulcei are expected to be produced as well Cheng and Feng (2022). Nevertheless, their detailed production mechanism is still unclear, where two different formalisms were adopted to describe the production of these states, i.e., the statistical thermal method Andronic et al. (2018); Botvina et al. (2023); Buyukcizmeci et al. (2018) and the coalescence model She et al. (2021); Zhou et al. (2025); Le Fèvre et al. (2019). Due to the differences in $N$-$\Lambda$ and NN-NN interactions, the production of hypernuclei has distinctive features in heavy-ion collisions compared with the corresponding normal nuclei of the same mass number Botta et al. (2012). Meanwhile, for heavy-ion collisions with beam energy close to hyperon threshold, higher-density nuclear matter is produced and the NN-NN and $N$-$\Lambda$ interactions as well as the mean field and kaon in-medium potentials are expected to impact the hyperon dynamics, e.g., collective flows, transverse momentum spectra, and rapidity distribution Yong (2025, 2024); Yong et al. (2025); Feng (2024); Wei et al. (2024); Zhou et al. (2025); Hartnack et al. (2012). As hypernuclei are produced in a fast expanding fireball created in heavy-ion collisions Reisdorf et al. (2010), the in-medium binding energy shift of hyperclusters maybe important for understanding hypernuclei production mechanism, particularly at beam energies around the hyperon threshold.
@@
In such cases, in this work we investigate the clustering phenomenon of $\Lambda$ hypernuclei in nuclear medium. In our previous work we have proposed a unified treatment for investigating the clustering phenomenon of normal nuclei inside a spherical Wigner-Seitz cell employing RMF models Xia (2025), which will be extended to the hyperonic sector here. The clusters are fixed by solving the Dirac equations imposing the Dirichlet-Neumann boundary condition Negele and Vautherin (1973), while the nuclear medium is treated with Thomas-Fermi approximation and assume constant densities. By calibrating the coupling constants $g_{\sigma\Lambda}$ and $g_{\omega\Lambda}$ according to the experimental separation energies of $\Lambda$ hyperons $B_{\Lambda}$ for single $\Lambda$ hypernuclei Hashimoto and Tamura (2006); Gal et al. (2016), the properties of $\Lambda$ hypernuclei in vacuum and in nuclear matter can be fixed.
@@ II Theoretical framework
The Lagrangian density for RMF models of static systems can be obtained with
@@
where $G_{n\kappa}(r)/r$ and $F_{n\kappa}(r)/r$ are the upper and lower components for the radial wave functions, $Y_{jm}^{l}(\theta,\phi)$ is the spinor spherical harmonics, $\kappa$ is connected to the angular momenta $(l,j)$ via $\kappa=(-1)^{j+l+1/2}(j+1/2)$. The Dirac equation is then reduced into
@@
where reflective boundary conditions are imposed at $r=0$ and the edge of Wigner-Seitz cell with $r=R_{\mathrm{W}}$ Xia et al. (2021). Note that the global charge neutrality is imposed by adding negatively charged electron background $n_{e}$ so that $A(R_{\mathrm{W}})=\nabla A(R_{\mathrm{W}})=0$, avoiding the possible divergence in Coulomb potential. Due to the reduction of Coulomb energy Xia (2025), this will lead to an increment of binding energies $B_{0}$ as indicated in Table 3 in comparison with that of vacuum $B$ in Table 2. For small systems such as heavy-ion collisions, the charge neutrality is not required and the variation of binding energies can be fixed by replacing $B_{0}$ with $B$. The source currents are comprised of contributions from both the nuclear medium (gas) and cluster, i.e.,
@@
The scalar ($n_{i}^{\mathrm{s}}$) and vector ($n_{i}$) densities of clusters (with given particle numbers $N_{i}$) are fixed according to the radial wave functions, i.e.,
@@
Note that the center-of-mass corrections for proton radius $\delta R_{p}^{2}$ are small and often neglected, for simplicity, in this work we neglect the mass difference between nucleons and hyperons in estimating $\delta R_{p}^{2}$ as the center-of-mass coordinate $\vec{R}_{\mathrm{c.m.}}=\sum_{i}^{A}m_{i}\vec{r}_{i}/\sum_{i}^{A}m_{i}$ varies little if we take $m_{p}=m_{n}=m_{\Lambda}$.
@@
with $g(x)=x(2x^{2}+1)\sqrt{x^{2}+1}-\mathrm{arcsh}(x)$. Here $\varepsilon_{ki}$ is the single-particle energy of particle $i$ in a cluster fixed by solving Dirac Eq. (8). The center-of-mass energy needs to be subtracted from Eq. (29) and can be fixed by Bender et al. (2000)
@@
where $S_{i,\mathrm{gas}}$ and $V_{i,\mathrm{gas}}$ are the scalar and vector mean-field potentials of baryon $i$ in the uniform nuclear medium. Note that taking the surrounding medium as a nuclear gas is a simplification of the real situation, where there may exist various clusters in the medium. In the gRDF approach with the binding energy shifts treated with a $\sigma$-cluster coupling, this could lead to a reduction of binding energies of clusters as $S_{i,\mathrm{gas}}$ is reduced in comparison with uniform nuclear gas scenarios at same $n_{i,\mathrm{gas}}$ Custódio et al. (2021).
@@ III Results and discussions
In this work we adopt the relativistic density functionals DD-LZ1 Wei et al. (2020), TM1 Sugahara and Toki (1994b), and PK1 Long et al. (2004), where the corresponding saturation properties of nuclear matter are listed in Table. 1. Note that the functionals PK1 and DD-LZ1 were widely adopted in our previous investigations on the properties of (hyper)nuclei Liu et al. (2018); Ding et al. (2023), while that of TM1 was extensively used for supernova simulations, i.e., the Shen EoSs Shen et al. (2011). Meanwhile, according to the state-of-the-art constraints on nuclear matter properties Russotto et al. (2016); Le Fèvre et al. (2016); Zhang et al. (2020); Essick et al. (2021); Huth et al. (2022), the functional DD-LZ1 Wei et al. (2020) predicts the most realistic equation of state.
@@ III.1 $\Lambda$ hypernuclei
We then investigate the properties of $\Lambda$ hypernuclei based on these effective NN-NN interactions in the absence of pairing interactions. The negligence of pairing interactions is justified for double magic nuclei since the formation of pairs is hindered by the large shell gap, while for open-shell nuclei this may lead to slight deviations of binding energies ($\sim 6/\sqrt{A}$ MeV) from the experimental values Zubov et al. (2009). For the $N$-$\Lambda$ interactions, we take the depth of $\Lambda$ mean field potential in symmetric nuclear matter (SNM) at saturation density to be
@@
which is expected to reproduce the single-$\Lambda$ binding energies irrespective of the particular values of $g_{\sigma\Lambda}$ and $g_{\omega\Lambda}$ Bouyssy and Hufner (1976); Millener et al. (1988); Wang et al. (2013); Sun et al. (2018); Rong et al. (2021). Consider the two-solar-mass constraint on neutron stars’ masses Sun et al. (2018, 2019), we take $g_{\sigma\Lambda}=g_{\sigma N}$ and then fix $g_{\omega\Lambda}$ according to Eq. (35), which gives $g_{\omega\Lambda}\approx 1.14g_{\omega N}$. Note that slightly smaller values for $g_{\sigma\Lambda}$ and $g_{\omega\Lambda}$ were adopted in our previous study Sun et al. (2018, 2019), which barely affect our predictions on single-$\Lambda$ binding energies in $\Lambda$ hypernuclei as long as Eq. (35) is satisfied.
@@
Figure 1: Binding energies of $\Lambda$ hyperon $B_{\Lambda}$ in the $1s_{1/2}$ state of hypernuclei predicted by the relativistic density functionals DD-LZ1 Wei et al. (2020), TM1 Sugahara and Toki (1994b), and PK1 Long et al. (2004). The corresponding experimental values are indicated by the open circles for comparison Hashimoto and Tamura (2006); Gal et al. (2016). The binding energy per nucleon $B_{\mathrm{core}}/(A-1)$ for the nuclear cores are presented as well. The element symbols (proton numbers $N_{p}$) are indicated for both $\Lambda$ hypernuclei and their core nuclei with $A\rightarrow A-1$, where each nucleus is presented independently.
@@
The binding energies of $\Lambda$ hyperon in the $1s_{1/2}$ states of $\Lambda$ hypernuclei ($^{4}_{\Lambda}$H, $^{4}_{\Lambda}$He, $^{7}_{\Lambda}$He, $^{7}_{\Lambda}$Be, $^{8}_{\Lambda}$Li, $^{10}_{\Lambda}$Be, $^{10}_{\Lambda}$B, $^{12}_{\Lambda}$B, $^{12}_{\Lambda}$C, $^{13}_{\Lambda}$C, $^{13}_{\Lambda}$C, $^{16}_{\Lambda}$N, $^{16}_{\Lambda}$O, $^{28}_{\Lambda}$Si, $^{32}_{\Lambda}$S, $^{40}_{\Lambda}$Ca, $^{51}_{\Lambda}$V, $^{52}_{\Lambda}$V, $^{89}_{\Lambda}$Y, $^{139}_{\Lambda}$La, and $^{208}_{\Lambda}$Pb) are then estimated with Eq. (1). The obtained results are presented in Fig. 1, while the corresponding experimental values are indicated as well for comparison Hashimoto and Tamura (2006); Gal et al. (2016). Evidently, the relativistic density functionals adopted in this work generally well reproduce the experimental binding energies of $\Lambda$ hyperon in $\Lambda$ hypernuclei in a large mass range of $A=4\sim 208$, where the corresponding root-mean-square deviations of $B_{\Lambda}$ from the experimental values are $\Delta=1.30$, 0.74, and 0.86 MeV when the functionals DD-LZ1 Wei et al. (2020), TM1 Sugahara and Toki (1994b), and PK1 Long et al. (2004) are adopted for the effective $NN$ interactions. The corresponding mean squared error are $\delta=8.54\%$, 9.85%, and 7.34%, so that the single $\Lambda$ binding energies are sufficiently well reproduced considering the uncertainties in the experimental data. Note that the deviations may be further reduced if we readjust the coupling constants and consider more realistic scenarios, e.g., pairing, deformation, and charge symmetry breaking Rong et al. (2025); Sun et al. (2025). The binding energy per nucleon $B_{\mathrm{core}}/(A-1)$ for the nuclear cores are presented as well. For light nuclei such as 3H and 3He, their binding energies per nucleon are larger than the binding energies of $\Lambda$ hyperon for $^{4}_{\Lambda}$H and $^{4}_{\Lambda}$He, which is mainly due to the weaker $N$-$\Lambda$ interaction than that of NN-NN interaction. As mass number ($A\gtrsim 8$) increases, nucleons eventually occupy single-particle states of larger energies, while $\Lambda$ hyperons still occupy the $1s_{1/2}$ state, so that the binding energies of $\Lambda$ hyperon eventually surpass that of nucleons.
@@
Table 2: Binding energies $B$ and charge radii $R_{\mathrm{ch}}$ for finite nuclei predicted by the relativistic density functional DD-LZ1 Wei et al. (2020). By adding one (two) $\Lambda$ hyperon(s) to these nuclei, the hyperon separation energies $B_{\Lambda}$ ($B_{\Lambda\Lambda}$) of $\Lambda$ ($\Lambda$$\Lambda$) hypernuclei as well as their charge radii $R_{\mathrm{ch}}^{\Lambda}$ ($R_{\mathrm{ch}}^{\Lambda\Lambda}$) are obtained and listed here. The experimental binding energies $B^{\mathrm{exp}}$ and charge radii $R_{\mathrm{ch}}^{\mathrm{exp}}$ for the core nuclei are indicated as well Angeli and Marinova (2013); Huang et al. (2021); Wang et al. (2021).
Core $B/A$ $B^{\mathrm{exp}}/A$ $R_{\mathrm{ch}}$ $R_{\mathrm{ch}}^{\mathrm{exp}}$ $B_{\Lambda}$ $R_{\mathrm{ch}}^{\Lambda}$ $B_{\Lambda\Lambda}/2$ $R_{\mathrm{ch}}^{\Lambda\Lambda}$
MeV MeV fm fm MeV fm MeV fm
3H 3.747 2.827 2.090 1.759 1.996 2.164 1.923 2.187
3He 3.419 2.573 2.337 1.966 2.064 2.407 1.972 2.414
4He 6.832 7.074 1.988 1.675 2.710 2.030 2.644 2.020
5Li 5.107 5.266 3.529 4.383 2.667 4.070 2.442
8Be 5.220 7.062 2.464 7.503 2.357 6.969 2.292
12C 7.663 7.680 2.354 2.470 11.48 2.313 11.05 2.292
16O 8.028 7.976 2.671 2.699 10.98 2.637 10.68 2.616
40Ca 8.610 8.551 3.440 3.478 16.64 3.415 16.38 3.397
48Ca 8.659 8.667 3.440 3.477 18.93 3.416 18.67 3.399
208Pb 7.871 7.867 5.485 5.501 25.49 5.468 25.35 5.452
[…]
@@
Figure 2: Density profiles of 4He, $^{5}_{\Lambda}$He, and $^{6}_{\Lambda\Lambda}$He in vacuum, where the relativistic density functional DD-LZ1 Wei et al. (2020) is employed for the effective NN-NN interactions and $g_{\omega\Lambda}=1.143g_{\omega N}$ ($g_{\sigma\Lambda}=g_{\sigma N}$) for the $\Lambda$-NN interactions.